OpenStudy (anonymous):

Your friend runs up to you, scared that he is not ready for the upcoming quadratics test. To help him study, you will create four different quadratic functions. Then demonstrate to him how to rewrite each function as a group of factors, if possible. The function f(x) is a difference of squares. The function g(x) is a sum of squares. The function h(x) is a perfect square trinomial. The function j(x) can only have a GCF factored out of it.

OpenStudy (anonymous):

I'm confused on the question itself

OpenStudy (anonymous):

How would I even demonstrate how to rewrite each function as a group of factors, if possible?

OpenStudy (anonymous):

@pooja195 @dan815 @Kainui

OpenStudy (anonymous):

@TheCleverOne ?

OpenStudy (anonymous):

@Serenity74 @fig-jig @Safa102

OpenStudy (dumbcow):

difference of squares: \[x^2 -a^2 = (x-a)(x+a)\] you can create any function by picking a num for "a" sum of squares: \[x^2 + a^2\] this cannot be factored into a group of factors perfect square: \[x^2 +2ax + a^2 = (x+a)^2\]

OpenStudy (anonymous):

thank you so much for helping first of all and okay so im not sure if i understand

OpenStudy (anonymous):

so just plug in a number to a in the first formula you gave ? @dumbcow

OpenStudy (dumbcow):

yes "a" can be any integer

OpenStudy (anonymous):

thanks @dumbcow